OpenStudy (anonymous):

The system of equations is coincident. What are the missing values? 3x+y=7 9x + ___ y= ______

OpenStudy (anonymous):

@mathsciencehistory

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (anonymous):

3x + y = 7 3(3x + y = 7) 9x + 3y = 21 --- the missing values are: 3 21

OpenStudy (anonymous):

help any

OpenStudy (anonymous):

yes it did thanks!